# Decoding the AI Pen: Techniques and Challenges in Detecting AI-Generated Text

### v1

@@ Abstract.
conference: ; ; copyright: none
@@ 1. Introduction
Leveraging AI-generated text could be regarded as an efficacious strategy for alleviating the aforementioned challenges. However, the convergence of AI-generated content with human-written texts has reached a point where discerning between the two has become increasingly intricate. The task of distinguishing LLM-generated text from human-written content presents a dual challenge.
@@
Figure 1. An overview of responsible AI-generated text study, with an emphasize on detection techniques and their challenges.
On one hand, identifying disparities can enhance the quality of AI-generated material. On the other hand, this endeavor also complicates the identification process. In recent years, researchers have proposed various methodologies for detecting AI-generated text . These efforts contribute to the ongoing exploration of this landscape.
This work aims to comprehensively investigate responsible AI-generated text. To this end, first, we highlight the risks and misuses associated with such text, while also discussing common mitigation strategies. As a significant solution, we thoroughly explore AI-generated text, concentrating on thematic categorization and assessing their constraints and vulnerabilities. Additionally, we conduct a theoretical exploration to assess the feasibility and potential of detecting AI-generated text. By adopting a theoretical lens, we seek to determine whether such detection is achievable or if detection remains an elusive goal within the domain of generative AI.
The rest of this paper is organized as follows: In section 2, we discuss the potential risks arising from the misuse of LLMs. Section 3 provides a comprehensive categorization of various text detection strategies, emphasizing their role as mitigation techniques. Subsequently, in Section 4, we describe the vulnerabilities inherent in these strategies. Moving forward, Section 5 engages in a theoretical exploration of the feasibility of detection, followed by section 6 where we propose new avenues of research. Finally, in section 7, we conclude. Figure 1 demonstrates an overview of the paper.
@@ 2. Risks and Misuse of AI-Generated Text
2.1. Discrimination, Toxicity and Harms
@@
Furthermore, the research reveal that specific genders and ethnicities are disproportionately susceptible to encountering toxic content. They posit that this phenomenon arises from an over-reliance on reinforcement learning with human feedback (RLHF) as a mechanism to mitigate model toxicity. However, the feedback provided to the model may carry inherent biases. For instance, feedback related to toxicity concerning different genders might be influenced by the representation of those genders.
@@
$2.$ Factual Inconsistency and Unreliability of AI Responses
@@
Significantly, the study highlights that users encounter challenges in assessing the accuracy of LLM-generated responses. Their judgment is influenced by the tone and style of the LLMs, which may introduce bias. This issue is extremely importance, as biases against uncertain text may impact the training and evaluation of LLMs.
To mitigate such mistakes, various strategies have been proposed through fine-tuning (Lewkowycz et al., 2022; Rajani et al., 2019; Zelikman et al., 2022), prompt engineering techniques such as verification, scratchpads  (Cobbe et al., 2021; Nye et al., 2022), Chain-of-Thought (CoT) (Wei et al., 2022), Reversing Chain-of-Thought (Xue et al., 2023) (RoCT), RLHF (Ziegler et al., 2019; Christiano et al., 2017), iterative self-reflection (Shinn et al., 2023; Madaan et al., 2023), self-consistency (Wang et al., 2023c), society of minds strategy (Du et al., 2023), pruning truthful datasets (Christiano et al., 2023), adjusting the system parameters to limit model creativity (Muneeswaran et al., 2023), external knowledge retrieval (Guu et al., 2020) and training-free methods with likelihood estimation (Kadavath et al., 2022).
@@ 2.3. Copyright Infringement and Plagiarism
LLMs could pose a substantial risk to academic writing by elevating the likelihood of copyright violations and plagiarism. For example, writers might employ LLMs to produce articles without original composition, while students could resort to LLMs for completing their assignments. These practices erode academic integrity and subvert the objectives of assignments and examinations (Khalil and Er, 2023; Stokel-Walker, 2022).
@@
Another crucial aspect to consider is the generalizability of AI-generated text detectors. These detectors should perform effectively on unseen data across diverse dimensions, including: models, languages and domains. Achieving robustness and reliability across these dimensions contributes to the overall effectiveness of machine-generated text detectors in real-world scenarios.
@@ 2.4. Misinformation Dissemination
However, research have shown that expanding the context size has minimal impact on mitigating the performance degradation that is caused by misinformation. Consequently, the straightforward strategy of diluting misinformation by increasing context size proves ineffective for defending against it (Pan et al., 2023b). A more straightforward approach involves prompting LLMs to issue warnings regarding misleading information. For instance, readers could be advised: “Be cautious, as some of the text may be intentionally misleading” .
@@ 3.1. Supervised Detection
For example, malicious actors have the potential to elude detection by leveraging their access to human.written texts present in the training set and detector rankings. Moreover, within a white-box context, attackers can undermine detector training—a concerning scenario, especially considering that numerous detectors rely on commonly used datasets, making them susceptible even to basic attacks. Additionally, these techniques are prone to the paraphrasing attack, where a paraphrased layer is added to the generative text model, allowing deception of any detector, including those utilizing supervised neural networks (Krishna et al., 2023; Sadasivan et al., 2023).
@@ 3.2. Zero-shot Detection
Another line of research employ pre-trained models as zero-shot classifiers to discern text written by AI, eliminating the necessity for supplementary training or data collection (Su et al., 2023; Zer, 2023; Wang et al., 2023a; Gehrmann et al., 2019). According to (Mitchell et al., 2023), commonly employed techniques often set a threshold for the predicted per-token log probability to identify AI-generated texts. This approach is grounded in the observation that passages generated by AI often exhibit a negative log probability curvature. While this method mitigates the risk of data poisoning and minimizes data and resource requirements, it remains vulnerable to other attacks such as spoofing (Shayegani et al., 2023) and paraphrasing (Krishna et al., 2023; Sadasivan et al., 2023).
$3.$ Retrieval-based Detection
An alternative avenue of research utilize information retrieval methods to differentiate between texts written by humans and those generated by machines. This is achieved by comparing a given text with a database of texts created by LLMs and identifying semantically similar matches (Krishna et al., 2023; Sadasivan et al., 2023).
Nevertheless, these approaches are impractical for real-world use due to their reliance on an extensive and up-to-date database of AI-generated texts. Such databases can be computationally costly or may not be available across all domains, tasks, or models. Additionally, like other techniques, these methods are susceptible to paraphrasing and spoofing attacks (Krishna et al., 2023; Sadasivan et al., 2023; Wolff, 2020; Liang et al., 2023).
Related papers Method Main Idea Vulnerabilities
(Solaiman et al., 2019; Bakhtin et al., 2019; Antoun et al., 2023; Zhan et al., 2023; Li et al., 2023) Supervised detection To fine-tune a model on sets of AI and human generated texts. Training on commonly used datasets, makes it vulnerable to most attacks including paraphrasing.
 (Su et al., 2023; Zer, 2023; Mitchell et al., 2023; Gehrmann et al., 2019; Wang et al., 2023a; Guo and Yu, 2023) Zero-shot detection To use a pre-trained language model in zero-shot settings. Reduces the risk of data poisoning attacks and eliminates data and resource over- heads, but it is still susceptible to other adversarial attacks like spoofing and paraphrasing.
 (Krishna et al., 2023; Sadasivan et al., 2023; Wolff, 2020; Liang et al., 2023) Retrieval-based detection Apply methods of information retrieval to match a given text with a collection of texts generated by LLMs and finding similarities in meaning. It is impractical because it requires a large and updated collection of texts, which is computationally expensive, or may be unavailable for all domains, tasks or models. It is also vulnerable to paraphrasing and spoofing attacks.
(Kirchenbauer et al., 2023a; Yang et al., 2023a; Kirchenbauer et al., 2023b; Sadasivan et al., 2023) Watermarking To use a model signature in the produced text outputs to stamp particular pattern. The most trustworthy strategy, but is shown to be fundamentally impossible for generative models. It is susceptible to attacks such as rewording and spoofing.
(Yu et al., 2023; Yang et al., 2023b; Mitchell et al., 2023; Su et al., 2023; Zhang et al., 2023) Feature-based detection To identify and classify based on extracted discriminating features. Susceptible to adversarial attacks such as paraphrasing.
@@ 3.4. Watermarking-based Detection
Another alternative avenue, referred to as watermarking techniques, employs a model signature within the generated text outputs to imprint specific patterns. For instance, Kirchenbauer et al. (Kirchenbauer et al., 2023a) propose a soft watermarking approach that categorizes tokens into green and red lists, facilitating the construction of these patterns. In this approach, a watermarked LLM selects a token, with a high likelihood, from the green list based on its preceding token. Remarkably, these watermarks often remain imperceptible to human observers.
To better understand the technique proposed by Kirchenbauer et al., assume an autoregressive language model is trained on a vocabulary $V$ of size $|V|$. Given a sequence of tokens as input at step $t$, a language model predicts the next token in the sequence by outputting a vector of logit scores $l_{t}\in R^{|V|}$ with one entry for each item in the vocabulary. A random number generator is seeded with a context window of $h$ preceding tokens, based on a pseudo-random function (PRF) $f:Nh\rightarrow N$. With this random seed, a subset of tokens of size $\gamma|V|$, where $\gamma in(0,1)$ is green list size, are “colored” green and denoted $G_{t}$. Now, the logit scores $l_{t}$ are modified such that witha hardness parameter $\sigma>0$:
@@
Despite previous indications, watermark-based techniques remain both theoretically and practically vulnerable to rewording attacks. Notably, even LLMs safeguarded by watermarking schemes are susceptible to spoofing attacks. In these attacks, human adversaries inject their own text into human-generated content, creating an illusion that the material originated from LLMs (Sadasivan et al., 2023). The interested reader is referred to (Sadasivan et al., 2023) for more details.
Moreover, unless all highly successful LLMs are uniformly safeguarded, watermarking remains an ineffective strategy for preventing LLM exploitation. Thus, the practical applicability of watermarking is restricted, especially in scenarios where only black-box LLMs are accessible. Due to API providers withholding probability distributions for commercial reasons, most third-party developers creating API-based applications cannot independently watermark text.
@@
They have observed that even after undergoing automated and human paraphrasing, watermarks can still be discerned. When a sufficient number of tokens are identified, paraphrased versions tend to inadvertently leak n-grams or even larger segments of the original text. Hence, high-confidence detection remains possible despite these attacks weakening the effectiveness of the watermark. The study proposes interpreting watermarking reliability as a function of text length. Surprisingly, even when attempting to remove the watermark intentionally, human writers struggle to do so if the text exceeds 1,000 words. This interpretation emerges as a significant characteristic of watermarking. Ultimately, according to this work, watermarking remains the most dependable strategy, as alternative paradigms like retrieval and loss-based detection have not demonstrated substantial improvements with increasing text length.
Despite prior findings, a recent study by Zhang et al. (Zhang et al., 2023) demonstrates that, under some assumptions, no robust watermarking scheme can prevent an attacker from removing the watermark without significantly degrading the output quality. We will further discuss this finding in section 5.
@@ 3.5. Detection via Discriminating Features
Where $\Omega$ is a set of $K$ samples of an LM decoder and $\hat{S}=LM(S_{1})$ and $S_{2}$ is the human written ground truth. “In both black-box and white-box scenarios, two parameters significantly impact detection accuracy: the truncation ratio and the number of re-prompting iterations $K$. This strategy reveals substantial discrepancies between AI-generated and human-written texts. This approach highlights substantial disparities between AI-generated and human-written texts.
Another distinguishing feature is the vulnerability of text to manipulations. Both AI-generated and human-written texts can be adversely affected by minor alterations, such as word replacements. However, recent research have revealed that machine-generated text is particularly prone to such manipulations (Mitchell et al., 2023; Su et al., 2023). For instance, Su et al. (Su et al., 2023) propose a metric called the Log-Likelihood Log-Rank Ratio (LRR) to quantify the sensitivity of LLMs to perturbations:
@@
where small perturbations are applied on the target text $x$ to produce the perturbed text $\tilde{x}_{p}$. This work demonstrates that the LRR tends to be larger for machine-generated text, making it a useful discriminator between machine-generated and human-written content. One plausible explanation is that, in machine-generated text, the log rank stands out more prominently than the log likelihood, resulting in a distinct pattern that LRR captures.
The rationale behind NPR lies in the fact that both machine-generated and human-written texts experience adverse effects from minor alterations. Specifically, the log rank score tends to increase after perturbations. However, machine-generated text is particularly vulnerable to such alterations, resulting in a more pronounced increase in the log rank score following perturbation. As a result, this pattern suggests a higher NPR score for AI-generated texts (Su et al., 2023).
@@ 4. Vulnerabilities of Detection Techniques
Despite various methods mentioned above, most categories of detection strategies are susceptible to paraphrasing or spoofing attacks. To tackle this, retrieval-based detectors have been employed as detection strategies. As mentioned earlier, these detectors store the output of LLMs in a database and perform semantic searches to extract the best matches. This approach enhances the detector’s resilience against paraphrasing attacks. However, there are privacy concerns related to storing user-LLM conversations. Also, this technique is ineffective when dealing with recursive paraphrasing (Sadasivan et al., 2023).
In addition, researchers have discovered that by effectively optimizing prompts, LLMs can successfully evade many detection techniques. For example, Lu et al. (Lu et al., 2023) propose a novel method called Substitution-based In-Context example Optimization (SICO), where discriminating features are extracted from human and AI-generated texts. These features, along with a paraphrasing prompt, are combined and fed to the LLM to modify its output. The prompt is carefully optimized through word and sentence replacements, aiming to minimize the chances of detection while maximizing the similarity between human and AI-generated texts. The results strongly demonstrate the vulnerability of existing methods.
Despite being considered an effective detection strategy, watermarking faces several challenges. Firstly, unless all LLMs are uniformly safeguarded, watermarking remains ineffective. Secondly, its practical applicability is limited, especially when dealing with black-box language models. Thirdly, API providers often withhold probability distributions, preventing third-party developers from independently watermarking text. Lastly, recent research suggest that no robust watermarking scheme can prevent attackers from removing watermarks without significantly degrading output quality. Therefore, watermarking generative models might be fundamentally unachievable, and alternative approaches are necessary to protect the intellectual property of model developers and LLM users. Table 1 illustrates an overview of detection strategies, highlighting the vulnerabilities associated with each category.
5. Possibility of Detection Through a Theoretical Lens
Given the increasing interest in detecting text generated by LLMs, researchers have recently investigated the theoretical aspects of this task. They explore the fundamental feasibility and limitations associated with identifying LLM-generated text.
Sadasivan et al. for instance, present a discovery of impossibility (Sadasivan et al., 2023). They assert that:as language models become increasingly sophisticated and adept at emulating human text, the effectiveness of even the best-possible detectors diminishes significantly. Sadasivan et al. establish an upper-bound for the area under the Receiver Operating Characteristic (ROC) curve of any decoder $\mathcal{D}$ as:

### v2

@@ Abstract.
LLM, AI-generated Text Detection; Responsible AI; Watermarking; Paraphrasing Attacks; Data Poisoning
journalyear: 2024copyright: rightsretainedconference: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 25–29, 2024; Barcelona, Spainbooktitle: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD ’24), August 25–29, 2024, Barcelona, Spaindoi: 10.1145/3637528.3671463isbn: 979-8-4007-0490-1/24/08ccs: Computing methodologies Natural language generationccs: Computing methodologies Neural networksccs: Security and privacy Social engineering attacks
@@ 1. Introduction
Figure 1. An overview of responsible AI-generated text study, with an emphasize on detection approaches and their challenges.
Leveraging AI-generated text could be regarded as an efficacious strategy for alleviating the aforementioned challenges. However, the convergence of AI-generated content with human-written texts has reached a point where discerning between the two has become increasingly intricate. The task of distinguishing LLM-generated text from human-written content presents a dual challenge. On one hand, identifying disparities can enhance the quality of AI-generated material. On the other hand, this endeavor also complicates the identification process. In recent years, researchers have proposed various methodologies for detecting AI-generated text . These efforts contribute to the ongoing exploration of this landscape.
This work aims to comprehensively investigate responsible AI-generated text. To this end, first, we highlight the risks and misuses associated with such text, while also discussing common mitigation strategies. As a significant solution, we thoroughly explore AI-generated text, concentrating on thematic categorization and assessing their constraints and vulnerabilities. Additionally, we conduct a theoretical exploration to assess the feasibility and potential of detecting AI-generated text. By adopting a theoretical lens, we seek to determine whether such detection is achievable or if detection remains an elusive goal within the generative AI domain.
The rest of this paper is organized as follows: In section 2, we discuss the potential risks arising from the misuse of LLMs. Section 3 provides a comprehensive categorization of various text detection strategies, emphasizing their role as mitigation techniques. Subsequently, in Section 4, we describe the vulnerabilities inherent in these strategies. Moving forward, Section 4.3 engages in a theoretical exploration of the feasibility of detection, followed by section 5 where we propose new avenues of research. Finally, in section 6, we conclude. Figure 1 demonstrates an overview of the paper.
@@ 2. Risks and Misuse of AI-Generated Text
2.1. Discrimination, Toxicity, and Harms
@@
Furthermore, the research reveal that specific genders and ethnicities are disproportionately susceptible to encountering toxic content. Deshpande et al. posit that this phenomenon arises from an over-reliance on reinforcement learning with human feedback (RLHF) as a mechanism to mitigate model toxicity. However, the feedback provided to the model may carry inherent biases. For instance, feedback related to toxicity concerning different genders might be influenced by the representation of those genders.
@@
2.2. Factual Inconsistency and Unreliability of AI Responses
@@
Significantly, the study highlights that users encounter challenges in assessing the accuracy of LLM-generated responses. Their judgment is influenced by the tone and style of the LLMs, which may introduce bias. This issue is extremely important, as biases against uncertain text may impact the LLM training and evaluation.
To mitigate such mistakes, various strategies have been proposed through fine-tuning (Lewkowycz et al., 2022; Rajani et al., 2019; Zelikman et al., 2022), prompt engineering techniques such as verification, scratchpads  (Cobbe et al., 2021; Nye et al., 2022), Chain-of-Thought (CoT) (Wei et al., 2022), Reversing Chain-of-Thought (RCoT)  (Xue et al., 2023), RLHF (Ziegler et al., 2019; Christiano et al., 2017), iterative self-reflection (Shinn et al., 2023; Madaan et al., 2023), self-consistency (Wang et al., 2023c), society of minds strategy (Du et al., 2023), pruning truthful datasets (Christiano et al., 2023), adjusting the system parameters to limit model creativity (Muneeswaran et al., 2023), external knowledge retrieval (Guu et al., 2020) and training-free methods with likelihood estimation (Kadavath et al., 2022).
@@ 2.3. Copyright Infringement and Plagiarism
LLMs could pose a substantial risk to academic writing by elevating the likelihood of copyright violations and plagiarism. For instance, writers might employ LLMs to produce articles without original composition, while students could resort to LLMs for completing homework assignments. These practices erode academic integrity and subvert the objectives of assignments and examinations (Khalil and Er, 2023; Stokel-Walker, 2022),
@@
Another crucial aspect to consider is the generalizability of AI-generated text detectors. These detectors should perform effectively on unseen data across diverse dimensions, including: models, languages and domains. Achieving robustness and reliability across these dimensions contributes to the overall effectiveness of AI-generated text detectors in real-world scenarios.
@@ 2.4. Misinformation Dissemination
However, research have shown that expanding the context size has minimal impact on mitigating the performance degradation caused by misinformation . Consequently, the straightforward strategy of diluting misinformation by increasing context size proves ineffective for defending against it (Pan et al., 2023b).
A more straightforward approach involves prompting LLMs to issue warnings regarding misleading information. For instance, readers could be advised: “Be cautious, as some of the text may be intentionally misleading” .
@@ 3.1. Supervised Detection
For example, malicious actors have the potential to elude detection by leveraging their access to human.written texts present in the training set and detector rankings. Moreover, within a white-box context, attackers can undermine detector training—a concerning scenario, especially considering that numerous detectors rely on commonly used datasets, making them susceptible even to basic attacks. Additionally, these techniques are prone to the paraphrasing attack, where a paraphrased layer is added to the generative text model, allowing deception of any detector, including those utilizing supervised neural networks (Krishna et al., 2023; Sadasivan et al., 2024).
@@ 3.2. Zero-shot Detection
Another line of research employ pre-trained models as zero-shot classifiers to discern text written by AI, eliminating the necessity for supplementary training or data collection (Su et al., 2023; Zer, 2023; Wang et al., 2023a; Gehrmann et al., 2019). According to (Mitchell et al., 2023), commonly employed techniques often set a threshold for the predicted per-token log probability to identify LLM-generated texts. This approach is grounded in the observation that passages generated by AI often exhibit a negative log probability curvature.
While this method mitigates the risk of data poisoning attacks and minimizes data and resource requirements, it remains vulnerable to other attacks such as spoofing (Shayegani et al., 2023) and paraphrasing (Krishna et al., 2023; Sadasivan et al., 2024).
3.3. Retrieval-based Detection
An alternative avenue of research utilize information retrieval methods to differentiate between texts written by humans and those generated by machines. This is achieved by comparing a given text with a database of texts created by LLMs and identifying semantically similar matches (Krishna et al., 2023; Sadasivan et al., 2024).
Nevertheless, these approaches are impractical for real-world use due to their reliance on an extensive and up-to-date database of AI-generated texts. Such databases can be computationally costly or may not be available across all domains, tasks, or models. Additionally, like other detection techniques we discussed earlier, these methods are susceptible to paraphrasing and spoofing attacks (Krishna et al., 2023; Sadasivan et al., 2024; Wolff, 2020; Liang et al., 2023).
@@ 3.4. Watermarking-based Detection
Another alternative avenue, referred to as watermarking techniques, employs a model signature within the generated text outputs to imprint specific patterns.
For instance, Kirchenbauer et al. (Kirchenbauer et al., 2023a) propose a soft watermarking approach that categorizes tokens into green and red lists, facilitating the construction of these patterns. In this approach, a watermarked LLM selects a token, with a high likelihood, from the green list based on its preceding token. Remarkably, these watermarks often remain imperceptible to human observers.
To better understand the technique proposed by Kirchenbauer et al., assume an autoregressive language model is trained on a vocabulary $V$ of size $|V|$. Given a sequence of tokens as input at step $t$, a language model predicts the next token in the sequence by outputting a vector of logit scores $l_{t}\in R^{|V|}$ with one entry for each item in the vocabulary. A random number generator is seeded with a context window of $h$ preceding tokens, based on a pseudo-random function (PRF) $f:Nh\rightarrow N$. With this random seed, a subset of tokens of size $\gamma|V|$, where $\gamma\in(0,1)$ is green list size, are “colored green” and denoted $G_{t}$. Now, the logit scores $l_{t}$ are modified such that with a hardness parameter $\sigma>0$:
@@
Despite previous indications, watermark-based techniques remain both theoretically and practically vulnerable to rewording attacks. Notably, even LLMs safeguarded by watermarking schemes are susceptible to spoofing attacks. In these attacks, human adversaries inject their own text into human-generated content, creating an illusion that the material originated from LLMs (Sadasivan et al., 2024). The interested reader is referred to (Sadasivan et al., 2024) for more details.
Moreover, unless all highly successful LLMs are uniformly safeguarded, watermarking remains an ineffective strategy for preventing LLM exploitation.
Furthermore, the practical applicability of watermarking is restricted, especially in scenarios where only black-box language models are accessible. Due to API providers withholding probability distributions for commercial reasons, most third-party developers creating API-based applications lack the capability to independently watermark text.
@@
They observe that even after undergoing automated and human paraphrasing, watermarks can still be discerned. When a sufficient number of tokens are identified, paraphrased versions tend to inadvertently leak n-grams or even larger segments of the original text. Hence, high-confidence detection remains possible despite these attacks weakening the effectiveness of the watermark. The study proposes interpreting watermarking reliability as a function of text length. Surprisingly, even when attempting to remove the watermark intentionally, human writers struggle to do so if the text exceeds 1,000 words. This interpretation emerges as a significant characteristic of watermarking. According to this work, watermarking remains the most dependable strategy, as alternative paradigms like retrieval and loss-based detections have not demonstrated substantial improvements with increasing text length.
Despite prior findings, a recent study by Zhang et al. (Zhang et al., 2023) demonstrates that, under some assumptions, no robust watermarking scheme can prevent an attacker from removing the watermark without significantly degrading the output quality. We will further discuss this finding in section 4.3.
@@ 3.5. Detection via Discriminating Features
Where $\Omega$ is a set of $K$ samples of an LM decoder and $\hat{S}=LM(S_{1})$ and $S_{2}$ is the human written ground truth. “In both black-box and white-box scenarios, two parameters significantly impact detection accuracy: the truncation ratio and the number of re-prompting iterations $K$. This strategy reveals substantial discrepancies between AI-generated and human-written texts. It highlights substantial disparities between AI-generated and human-written texts.
Another distinguishing feature is the vulnerability of text to manipulations. Both AI-generated and human-written texts can be adversely affected by minor alterations, such as word replacements. However, recent research have revealed that AI-generated text is particularly prone to such manipulations (Mitchell et al., 2023; Su et al., 2023).
For instance, Su et al. (Su et al., 2023) propose a metric called the Log-Likelihood Log-Rank Ratio (LRR) to quantify the sensitivity of LLMs to perturbations:
@@
where small perturbations are applied on the target text $x$ to produce the perturbed text $\tilde{x}_{p}$.
This work demonstrates that the LRR tends to be larger for AI-generated text, making it a useful discriminator between AI and human-generated content. One plausible explanation is that, in AI-generated text, the log rank stands out more prominently than the log likelihood, resulting in a distinct pattern that LRR captures.
The rationale behind NPR lies in the fact that both AI-generated and human-written texts experience adverse effects from minor alterations. Specifically, the log rank score tends to increase after perturbations. However, AI-generated text is particularly vulnerable to such alterations, resulting in a more pronounced increase in the log rank score following perturbation. As a result, this pattern suggests a higher NPR score for AI-generated texts (Su et al., 2023).
Table 6 illustrates an overview of detection strategies, highlighting the vulnerabilities associated with each category.
@@ 4. Vulnerabilities of Detection Techniques
Despite various methods mentioned above, most categories of detection strategies are susceptible to paraphrasing or spoofing attacks. To tackle this, retrieval-based detectors have been employed as detection strategies but these detectors store the output of LLMs in a database and perform semantic searches to extract the best matches. Although it enhances the detector’s resilience against paraphrasing attacks, there are privacy concerns related to storing user-LLM conversations. Additionally, this technique is ineffective when dealing with recursive paraphrasing (Sadasivan et al., 2024). Moreover, researchers have discovered that by effectively optimizing prompts, LLMs can successfully evade many detection techniques.
For example, in a recent work (Lu et al., 2023), Lu et al. propose a novel method called Substitution-based In-Context example Optimization (SICO). In SICO, discriminating features are extracted from both human and AI-generated texts. These features, along with a paraphrasing prompt, are combined and fed to the LLM to modify its output. The prompt is carefully optimized through word and sentence replacements, aiming to minimize the chances of detection while maximizing the similarity between human and AI-generated texts. Evaluation results strongly demonstrate the vulnerability of existing methods.
Despite being considered an effective detection strategy, watermarking faces several challenges. Firstly, unless all LLMs are uniformly safeguarded, watermarking remains ineffective. Secondly, its practical applicability is limited, especially when dealing with black-box language models. Thirdly, API providers often withhold probability distributions, preventing third-party developers from independently watermarking text. Lastly, recent research suggest that no robust watermarking scheme can prevent attackers from removing watermarks without significantly degrading output quality. Therefore, watermarking generative models might be fundamentally unachievable, and alternative approaches are necessary to protect the intellectual property of model developers and LLM users. Figure 2 presents a concise overview of vulnerabilities in detection techniques.
Given the increasing interest in detecting text generated by LLMs, researchers have recently investigated the theoretical aspects of this task. They explore the fundamental feasibility and limitations associated with identifying LLM-generated text. In this section we explore some of the notable research studies on this topic.
Figure 2. A summary of detection vulnerabilities.
4.1. Impossibility via AUROC Upper-bound
In a new study, Sadasivan et al. present a discovery of impossibility (Sadasivan et al., 2024). They claim that:
As language models become increasingly sophisticated and adept at emulating human text, the effectiveness of even the best-possible detectors diminishes significantly.
Sadasivan et al. establish an upper-bound for the area under the Receiver Operating Characteristic (ROC) curve of any decoder $\mathcal{D}$: